in a different experiment it is observed that 252J of heat must be absorbed to raise the temp of 50.0g of Ni(s) from 20.0 C to 31.4 C
calculate the specific heat of Ni in both J/g Degree C and cal/g Degree C
1
Expert's answer
2015-11-06T04:12:07-0500
The specific heat of Ni is defined:
C = Q/(m∆T), where Q – the absorbed heat, m – the mass and ∆T – the change of temperature.
C = 252 J/(50.0 g × [31.4 oC – 20.0 oC]) = 0.442 J/(g· oC)
1 calorie equal 4.18 J, therefore Q = 252 J = 252 J/4.18 = 60.29 cal.
Hence, C = 60.29 cal/(50.0 g × [31.4 oC – 20.0 oC]) = 0.106 cal/(g· oC)